OpenStudy (anonymous):

anyone wanna help me on this question? Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)).

OpenStudy (anonymous):

and these are the answer choices if it helps: -40x + 70 -40x + 64 -40x + 24 -40x + 31

OpenStudy (ivettef365):

just substitute the x from f(x) for 10x-6

OpenStudy (anonymous):

like f(10x-6)= -4x + 7 ?

OpenStudy (ivettef365):

can you solve it like that

OpenStudy (ivettef365):

I meant -4x+7= -4(10x-6)+7

OpenStudy (anonymous):

oh okay. let me see what I get.

OpenStudy (ivettef365):

let me know what you got

OpenStudy (anonymous):

-40x + 31?

OpenStudy (anonymous):

Yeah, the answer is correct.